OpenStudy (anonymous):

What is b = 1 ----- ---- 15.7 2.5

OpenStudy (anonymous):

2.5 is 25, sorry.

OpenStudy (anonymous):

i do not get your question...

Directrix (directrix):

b/15.7 = 1/25 Cross Multiply (Means-Extremes Property) 25 b = 1*15.7 25b = 15.7 b = ? --> @Turtl_Johnson Check my arithmetic. Post your value for b and somebody will stop by to check it. Okay?